# Sacramento, CA, USA

### 10 Feb 2024 v2

@@ 3.5 Dataset
Field Value
financebench_id financebench_id_00859
doc_name VERIZON_2021_10K
doc_link https://www.verizon.com/about/sites/default/files/2021-Annual-Report-on-Form-10-K.pdf
question_type ’novel-generated’
question Among all of the derivative instruments that Verizon used to manage the exposure to fluctuations of foreign currencies exchange rates or interest rates, which one had the highest notional value in FY 2021?
answer Cross currency swaps. Its notional value was \$32,502 million.,
evidence_text Derivative Instruments We enter into derivative transactions primarily to manage our exposure to fluctuations in foreign currency exchange rates and interest rates. We employ risk management strategies, which may include the use of a variety of derivatives including interest rate swaps, cross currency swaps, forward starting interest rate swaps, treasury rate locks, interest rate caps, swaptions and foreign exchange forwards. We do not hold derivatives for trading purposes. The following table sets forth the notional amounts of our outstanding derivative instruments: (dollars in millions) At December 31, 2021 2020 Interest rate swaps \$ 19,779 \$ 17,768 Cross currency swaps 32,502 26,288 Forward starting interest rate swaps 1,000 2,000 Foreign exchange forwards 932 1,405
page_number 85

### 16 Mar 2024 v3

@@ Abstract
Keywords:
Retrieval Augmented Generation Document Chunking Document Pre-Processing Financial Domain Large Language Models
@@ 3.5 Dataset
Field
Value
financebench_id
financebench_id_00859
doc_name
VERIZON_2021_10K
doc_link
https://www.verizon.com/about/sites/default/files/2021-Annual-Report-on-Form-10-K.pdf
question_type
’novel-generated’
question
Among all of the derivative instruments that Verizon used to manage the exposure to fluctuations of foreign currencies exchange rates or interest rates, which one had the highest notional value in FY 2021?
answer
Cross currency swaps. Its notional value was \$32,502 million.,
evidence_text
Derivative Instruments We enter into derivative transactions primarily to manage our exposure to fluctuations in foreign currency exchange rates and interest rates. We employ risk management strategies, which may include the use of a variety of derivatives including interest rate swaps, cross currency swaps, forward starting interest rate swaps, treasury rate locks, interest rate caps, swaptions and foreign exchange forwards. We do not hold derivatives for trading purposes. The following table sets forth the notional amounts of our outstanding derivative instruments: (dollars in millions) At December 31, 2021 2020 Interest rate swaps \$ 19,779 \$ 17,768 Cross currency swaps 32,502 26,288 Forward starting interest rate swaps 1,000 2,000 Foreign exchange forwards 932 1,405
page_number
85